# Trajectory-Based Multi-Objective Hyperparameter Optimization for Model Retraining

### v1

@@ 4 Gaussian Process for Trajectory Prediction
where $K(\bm{z},Z)=K(Z,\bm{z})^{T}\in\mathbb{R}^{n}$ with $\left[K(\bm{z},Z)\right]_{i}=K(\bm{z},\bm{z}_{i})$. Refer to [33] for a comprehensive review of GPs. In each iteration of the vanilla BO method, a new input $\bm{z}^{\prime}\in\mathbb{Z}$ is selected by optimizing an acquisition function derived from the predictive mean $\mu$ and uncertainty $\Sigma$. Upon observing $y^{\prime}$ at $\bm{z}^{\prime}$, BO advances to the next iteration with the updated input and observation sets $Z=Z\cup\{\bm{z}^{\prime}\}$ and $Y=Y\cup\{y^{\prime}\}$.
As each input $\bm{z}=(\bm{x},t)$, we define the kernel function $K\left((\bm{x},t),(\bm{x}^{\prime},t^{\prime})\right)$ as the product of a standard kernel $K(\bm{x},\bm{x}^{\prime})$ over hyperparameter setting and a temporal kernel $K(t,t^{\prime})$ over epochs, with the latter capturing relationships across different epochs for a fixed hyperparameter setting. For instance, Swersky et al. [38] proposed an exponential decaying kernel to account for the validation loss that exponentially decreases over $t$. Belakaria et al. [3] used a linear kernel when modeling learning curves related to training costs. Given that learning curves for different performance measures may exhibit different characteristics, in this study we employ specific temporal kernels if their behavior is known a priori. By fitting a GP model for each objective function $f_{i}$, $i\in\{1,\dots,k\}$, we can predict the trajectory of any hyperparameter setting $\bm{x}$ and further improve the accuracy of its trajectory prediction by continuously monitoring the changes in model performance during iterative learning.
@@ 5 Trajectory-Based Bayesian Optimization Approach
Now we introduce an epoch-efficient algorithm named Trajectory-based Multi-Objective Bayesian Optimization (TMOBO) for solving the EMOHPO as defined in (1). This algorithm is particularly designed to efficiently navigate the trade-off model performances across multiple epochs by leveraging the insights obtained from trajectories. At its core, TMOBO features a trajectory-based acquisition function to sample hyperparameter settings and a trajectory-based early stopping mechanism to determine the number of epochs to train with each hyperparameter setting. The pseudo-code of TMOBO is presented in Algorithm 1.
@@
1:while the computational budget has not been exceeded do
2:     Fit $k$ GP models with $\bm{\mu}$ and $\bm{\Sigma}$ based on sets $Z$ and $Y$. $\triangleright$ refer to Section 4
3:     Sample a new $\bm{x}^{\prime}$ by maximizing the TEHVI acquisition function. $\triangleright$ refer to Section 5.1
@@
8:         Fit $k$ GP models with $\bm{\mu}$ and $\bm{\Sigma}$ based on sets $Z\cup Z^{\prime}$ and $Y\cup Y^{\prime}$.
9:         if EarlyStopping$(\bm{x}^{\prime},t^{\prime},\bm{\mu},\bm{\Sigma},F)$ is triggered then $\triangleright$ refer to Section 5.2
@@
13:     Augment $Z^{\prime}$ and $Y^{\prime}$ into $Z$ and $Y$ respectively. $\triangleright$ refer to Section 5.2
@@
Each iteration of TMOBO is centered around a two-step sampling strategy where a hyperparameter setting $\bm{x}^{\prime}$ and its corresponding number of epochs $t^{\prime}$ are determined successively. The iteration starts with fitting $\bm{\mu}=\left[\mu_{1},\dots,\mu_{k}\right]$ and $\bm{\Sigma}=\left[\Sigma_{1},\dots,\Sigma_{k}\right]$, with $\mu_{i}$ and $\Sigma_{i}$ representing the predictive mean and uncertainty for the $i$-th objective function. An unvisited setting $\bm{x}^{\prime}$ is then selected by maximizing an acquisition function that measures the potential improvement made by the trajectory of a hyperparameter setting as if it were to be fully trained. Thereafter, we start to train the ML model with $\bm{x}^{\prime}$, monitor the model performance, and predict its future trajectory epoch by epoch. Once the criterion for early stopping is met, the training for $\bm{x}^{\prime}$ is terminated to conserve the computational budget. Finally, TMOBO augments the primary datasets $Z$ and $Y$ by selecting the most informative observations associated with $\bm{x}^{\prime}$.
5.1 Trajectory-Based Expected Hypervolume Improvement
As per Definition 3, Hypervolume (HV) evaluates the quality of a set of solutions in the objective space without any prior knowledge of the actual Pareto-optimal front. The maximization over HV can yield a set of solutions that are converged to and well-distributed along the Pareto-optimal front, which makes HV one of the most popular indicators used in multi-objective optimization. As per Definition 4, Hypervolume Improvement (HVI) is built upon HV and quantifies the increase in HV as the gain brought by a solution.
@@ 5.1 Trajectory-Based Expected Hypervolume Improvement
Considering that the objective values of any out-of-sample $\bm{z}$ are unknown ahead of time, the direct computation of HVI for $\bm{z}$ becomes infeasible. Therefore, Expected Hypervolume Improvement (EHVI) [48, 8] has been used in the BO framework to estimate the gain of $\bm{z}$ by taking the expectation of HVI over the predictive distribution of its objective values, i.e.,
@@
Recall that in our study each $\bm{z}$ is composed of a hyperparameter setting $\bm{x}$ and a specific epoch number $t$. Obviously, EHVI determines a hyperparameter setting $\bm{x}$ purely based on its model performance after $t$ epochs without considering any valuable insights from the past observed or future potential model performance on the trajectory $Trj(\bm{x})$. Meanwhile, Due to its joint sampling of hyperparameter setting and training epochs, EHVI treats $\bm{z}_{1}=(\bm{x},t_{1})$ and $\bm{z}_{2}=(\bm{x},t_{2})$ as different pairs. This can lead to inefficiencies in iterative learning, as it allows for repeated evaluation on the same hyperparameter setting.
Theorem 1.
@@
Define $X_{Trj}^{*}:=\{\bm{x}\mid\bm{x}\in\mathbb{X},\exists t\in\mathbb{T},\nexists(\bm{x}^{\prime},t^{\prime})\in\mathbb{X}\times\mathbb{T},\bm{f}(\bm{x}^{\prime},t^{\prime})\prec\bm{f}(\bm{x},t)\}$. The Pareto-optimal set of $X_{Trj}^{*}\times\mathbb{T}$ is equal to the Pareto-optimal set of EMOHPO in (1).
@@
The theorem (see proof in Appendix A) above inspires the idea of identifying the set $X_{Trj}^{*}$ instead of $Z^{*}$, which allows us to search in a lower-dimensional space and to determine $\bm{x}$ regardless of $t$. Moreover, $X_{Trj}^{*}$ is in fact the set of all hyperparameter settings whose trajectories are the “best” or, in other words, contribute to the Pareto optimality. Therefore, in the BO framework, it can be approached by sequentially finding the trajectory that makes the most significant improvement. To this end, we introduce the Trajectory-based EHVI (TEHVI) that operates over $\bm{x}$ only and wraps $t$ into the trajectory $Trj(\bm{x})$,
$TEHVI(\bm{x}\mid F,\bm{r}):=\mathbb{E}\left[HVI\left(Trj(\bm{x})\mid F,\bm{r}\right)\right]=\mathbb{E}\left[HVI\left(\Big{\{}\bm{f}(\bm{x},t)\Big{\}}_{t=1}^{t_{max}}\mid F,\bm{r}\right)\right].$ (5)
@@
By maximizing TEHVI across the hyperparameter space, we attempt to locate the hyperparameter setting that has the best trajectory regarding the current front $F$. However, it is worth noting that TEHVI is equivalent to the joint EHVI of multiple positions along a trajectory, which has no known analytical form and becomes particularly complicated when $t_{max}$ is large. Therefore, following the previous studies on the fast computation of EHVI [13, 8], we resort to the Monte Carlo (MC) integration for approximating TEHVI, i.e.,
$TEHVI(\bm{x}\mid F,\bm{r})\approx\frac{1}{M}\sum_{m=1}^{M}HVI\left(\widehat{Trj}_{m}(\bm{x})\mid F,\bm{r}\right),$ (6)
@@
where $\widehat{Tr}_{m}(\bm{x}):=\{\widehat{\bm{f}}_{m}(\bm{x},t)\}_{t=1}^{t_{max}}$ denotes a predictive trajectory of $\bm{x}$ sampled from the joint posterior of GPs and $M$ denotes the total number of samples. To further alleviate the computational burden, we adopt the candidate search strategy that maximizes the approximated TEHVI over a fixed-size set of candidate hyperparameter settings. Each candidate is generated by adding a Gaussian perturbation to the evaluated hyperparameter setting whose trajectory has contributed to the current front the most. It has been shown that such a candidate search guarantees asymptotic convergence to the global optimum [34, 43]. More importantly, it enables the simultaneous computation of TEHVI for multiple candidates in a batch to significantly improve efficiency. Please refer to Appendix B for more details.
5.2 Early Stopping and Data Augmentation
@@
$t^{*}=\sup\big{\{}t\in\mathbb{T}\mid\bm{\mu}(\bm{x}^{\prime},t)-\beta^{1/2}\bm{\Sigma}(\bm{x}^{\prime},t)\prec\bm{y},\exists\bm{y}\in F\big{\}}.$ (7)
where $\beta$ is a predetermined constant controlling the confidence level. Inspired by the Lower Confidence Bound (LCB) [36], the conservative stopping epoch is the maximum number of epochs after which any future model performance is unlikely to improve the current front with a high probability. As the iterative learning procedure proceeds by epoch, the trajectory prediction is progressively updated based on the new observations on $Trj(\bm{x}^{\prime})$ and so is the conservative stopping epoch. The iterative learning procedure terminates once the current training epoch $t^{\prime}$ exceeds the conservative stopping epoch $t^{*}$.
Consequently, at the end of each iteration of TMOBO, we obtain the temporary datasets $Z^{\prime}=\{(\bm{x}^{\prime},t)\}_{t=1}^{t^{\prime}}$ and $Y^{\prime}=\{\bm{y}(\bm{x}^{\prime},t)\}_{t=1}^{t^{\prime}}$, where the actual stopping epoch $t^{\prime}$ can take any value from $\{1,\dots,t_{max}\}$. However, retaining all new observations in $Z^{\prime}$ and $Y^{\prime}$ for training GP models in subsequent iterations is inefficient, especially when $t_{max}$ is large. Following the active data augmentation used in [29, 3], we opt to augment only a subset of $Z^{\prime}$ to the primary dataset $Z$ by sequentially selecting an input $(\bm{x}^{\prime},t)\in Z^{\prime}$ at which the GP predictive uncertainty is highest. As more than one GP models are used to approximate the objectives, similar to the scenario considered in [3], the model uncertainty of $(\bm{x}^{\prime},t)$ is computed as the sum of normalized variances predicted by each GP model at $(\bm{x}^{\prime},t)$. Through this method, we can effectively control the increase in the size of the training set while ensuring the accuracy and training efficiency of GP models.
@@ 6 Numerical Experiments
In this section, we conduct a comprehensive empirical analysis of the performance of TMOBO on several synthetic and real-world benchmark problems that are formulated as EMOHPO. As demonstrated in Section 2, EMOHPO is essentially a multi-objective optimization problem defined over the combined space of hyperparameter settings and training epochs. Therefore, we compare TMOBO with several state-of-the-art MOBO methods in the literature, namely ParEGO [23], $q$EHVI [8], and $q$NEHVI [9]. To ensure a fair comparison, we enhance these MOBO methods by collecting all the observations $\{\bm{y}(\bm{x},1),\dots,\bm{y}(\bm{x},t)\}$ into their results when sampling at the pair $(\bm{x},t)$. For clarity, in the following discussion, we use ParEGO-T to represent the enhanced version of ParEGO, and similarly for $q$EHVI-T and $q$NEHVI-T.
Considering that all algorithms are stochastic, we perform 20 independent trial runs of each algorithm on each test problem. When measuring the performance of an algorithm, we compute the logarithm of the HV difference between the front found by the algorithm and the true Pareto-optimal front [8, 9]. As the true Pareto-optimal front is generally unknown, it is estimated by the best front identified from the collection of actual observations across all algorithms and trials. Similarly, the reference point is set as the least favorable solution with each dimension corresponding to the worst observed value of an objective so that it upper bounds all observations in the objective space for HV computation. See Appendix C.1 for more details on experimental settings.
6.1 Experiments on Synthetic Simulations

### v2

@@ 4 Gaussian Process for Trajectory Prediction
where $K(\bm{z},Z)=K(Z,\bm{z})^{T}\in\mathbb{R}^{n}$ with $\left[K(\bm{z},Z)\right]_{i}=K(\bm{z},\bm{z}_{i})$. Refer to [Rasmussen and Williams, 2005] for a comprehensive review of GPs. In each iteration of the vanilla BO method, a new input $\bm{z}^{\prime}\in\mathbb{Z}$ is selected by optimizing an acquisition function derived from the predictive mean $\mu$ and uncertainty $\Sigma$. Upon observing $y^{\prime}$ at $\bm{z}^{\prime}$, BO advances to the next iteration with the updated input and observation sets $Z=Z\cup\{\bm{z}^{\prime}\}$ and $Y=Y\cup\{y^{\prime}\}$.
As each input $\bm{z}=(\bm{x},t)$, we define the kernel function $K\left((\bm{x},t),(\bm{x}^{\prime},t^{\prime})\right)$ as the product of a standard kernel $K(\bm{x},\bm{x}^{\prime})$ over hyperparameter setting and a temporal kernel $K(t,t^{\prime})$ over epochs, with the latter capturing relationships across different epochs for a fixed hyperparameter setting. For instance, Swersky et al. [2014] proposed an exponential decaying kernel to account for the validation loss that exponentially decreases over $t$. Belakaria et al. [2023] used a linear kernel when modeling learning curves related to training costs. Given that learning curves for different performance measures may exhibit different characteristics, in this study we employ specific temporal kernels if their behavior is known a prior. See Appendix B.1 for an illustrative example of GP prediction. By fitting a GP model for each objective function $f_{i}$, $i\in\{1,\dots,k\}$, we can predict the trajectory $Trj(\bm{x})$ for any hyperparameter setting $\bm{x}$ and further improve the accuracy of its trajectory prediction by continuously monitoring the changes in model performance during iterative learning.
@@ 5 Trajectory-Based Bayesian Optimization Approach
Now we introduce an epoch-efficient algorithm named Trajectory-based Multi-Objective Bayesian Optimization (i.e., TMOBO) for solving the EMOHPO as defined in (1). This algorithm is particularly designed to efficiently navigate the trade-off model performances across multiple epochs by leveraging the insights obtained from trajectories. At its core, TMOBO features a trajectory-based acquisition function to sample hyperparameter settings and a trajectory-based early stopping mechanism to determine the number of epochs to train with each hyperparameter setting. The pseudo-code of TMOBO is presented in Algorithm 1.
@@
1:while computational budget has not been exceeded do
2:     Fit $k$ GPs with $\bm{\mu}$ and $\bm{\Sigma}$ based on sets $Z$ and $Y$.
@@
8:         Fit $k$ GPs with $\bm{\mu}$ and $\bm{\Sigma}$ based on sets $Z\cup Z^{\prime}$ and $Y\cup Y^{\prime}$.
9:         if EarlyStopping$(\bm{x}^{\prime},t^{\prime},\bm{\mu},\bm{\Sigma},F)$ triggered then
@@
13:     Augment $Z^{\prime}$ and $Y^{\prime}$ into $Z$ and $Y$ respectively.
@@
Each iteration of TMOBO is centered around a two-level sampling strategy where a hyperparameter setting $\bm{x}^{\prime}$ and its corresponding number of epochs $t^{\prime}$ are determined successively. The iteration starts with fitting $\bm{\mu}=\left[\mu_{1},\dots,\mu_{k}\right]$ and $\bm{\Sigma}=\left[\Sigma_{1},\dots,\Sigma_{k}\right]$, with $\mu_{i}$ and $\Sigma_{i}$ representing the predictive mean and uncertainty for the $i$-th objective function. An unvisited setting $\bm{x}^{\prime}$ is then selected by maximizing an acquisition function that measures the potential improvement made by the trajectory of a hyperparameter setting as if it were to be fully trained. Thereafter, we train the ML model with $\bm{x}^{\prime}$, monitor the model performance, and predict its future trajectory epoch by epoch. Once the criterion for early stopping is met, the training for $\bm{x}^{\prime}$ is terminated to conserve the computational budget. Finally, TMOBO augments the primary datasets $Z$ and $Y$ by selecting the most informative observations associated with $\bm{x}^{\prime}$.
5.1 Trajectory-Based Acquisition Function
As per Definition 3, Hypervolume (HV) evaluates the quality of a set of solutions in the objective space without any prior knowledge of actual Pareto-optimal front. The maximization over HV yields a set of solutions that are converged to and well-distributed along the Pareto-optimal front, which makes HV one of the most popular indicators used in multi-objective optimization. As per Definition 4, Hypervolume Improvement (HVI) is built upon HV and quantifies the increase in HV as the gain brought by a solution.
@@ 5.1 Trajectory-Based Acquisition Function
Given that the objective values of any out-of-sample $\bm{z}$ are unknown ahead of time, the direct computation of HVI for $\bm{z}$ is infeasible. Therefore, Expected Hypervolume Improvement (EHVI) [Zitzler et al., 2007, Daulton et al., 2020] has been used in the BO framework to estimate the gain of $\bm{z}$ by taking the expectation of HVI over the predictive distribution of its objective values, i.e.,
$EHVI(\bm{z}\mid F,\bm{r})$ $=\mathbb{E}\left[HVI(\bm{f}(\bm{z})\mid F,\bm{r})\right]$
@@
Recall that in our study each $\bm{z}$ is composed of a hyperparameter setting $\bm{x}$ and a specific epoch number $t$. Obviously, EHVI determines a hyperparameter setting $\bm{x}$ purely based on its model performance after $t$ epochs without considering any valuable insights from the past observed or future potential model performance on the trajectory $Trj(\bm{x})$. Meanwhile, due to its joint sampling of hyperparameter setting and training epochs, EHVI ignores the fact that model performances at $\bm{z}_{1}=(\bm{x},t_{1})$ and $\bm{z}_{2}=(\bm{x},t_{2})$ can be observed in a single query in EMOHPO. In order words, this means that redundant model training with the same hyperparameter setting is allowed, which can cause inefficiencies in the optimization of EMOHPO.
Lemma 1.
@@
Let $X_{Trj}^{*}$ denote the set of hyperparameter settings that belong to the Pareto-optimal set of EMOHPO, i.e., $X_{Trj}^{*}:=\{\bm{x}\in\mathbb{X}\mid\exists t\in\mathbb{T},\nexists(\bm{x}^{\prime},t^{\prime})\in\mathbb{X}\times\mathbb{T},\bm{f}(\bm{x}^{\prime},t^{\prime})\prec\bm{f}(\bm{x},t)\}$. Then, the Pareto-optimal set of EMOHPO is equivalent to the Pareto-optimal set of $X_{Trj}^{*}\times\mathbb{T}$.
@@
The above lemma (see proof in Appendix A) inspires the idea of solving EMOHPO by identifying the set $X_{Trj}^{*}$, which consists of hyperparameter settings whose trajectories are “best”, or more precisely, contribute to Pareto optimality, rather than directly searching for $Z^{*}$. By focusing on $X_{Trj}^{*}$ and subsequently observing the trajectory of each $\bm{x}^{*}\in X_{Trj}^{*}$, we effectively reduce the search space to a lower-dimensional domain, allowing us to determine $\bm{x}$ independently of $t$. Within the BO framework, it can be achieved by iteratively finding the trajectory that makes the most significant improvement. To this end, we introduce the Trajectory-based EHVI (TEHVI) that operates over $\bm{x}$ only and wraps $t$ into the trajectory $Trj(\bm{x})$,
$TEHVI(\bm{x}\mid F,\bm{r})$ $:=\mathbb{E}\left[HVI\left(Trj(\bm{x})\mid F,\bm{r}\right)\right]$
$=\mathbb{E}\left[HVI\left(\Big{\{}\bm{f}(\bm{x},t)\Big{\}}_{t=1}^{t_{max}}\mid F,\bm{r}\right)\right].$
@@
By maximizing TEHVI across the hyperparameter space, we attempt to locate the hyperparameter setting that has the best trajectory regarding the current front $F$. However, it is worth noting that TEHVI is equivalent to the joint EHVI of multiple positions along a trajectory, which has no known analytical form and becomes particularly complicated when $t_{max}$ is large. Therefore, following the previous studies on the fast computation of EHVI [Emmerich et al., 2006, Daulton et al., 2020], we resort to the Monte Carlo (MC) integration for approximating TEHVI, i.e.,
$TEHVI(\bm{x}\mid F,\bm{r})\approx\frac{1}{M}\sum_{m=1}^{M}HVI\left(\widehat{Trj}_{m}(\bm{x})\mid F,\bm{r}\right),$
@@
where $\widehat{Tr}_{m}(\bm{x}):=\{\widehat{\bm{f}}_{m}(\bm{x},t)\}_{t=1}^{t_{max}}$ denotes a predictive trajectory of $\bm{x}$ sampled from the joint posterior of GPs and $M$ denotes the total number of samples. To further alleviate the computational burden, we adopt the candidate search strategy that maximizes the approximated TEHVI over a fixed-size set of candidate hyperparameter settings. Each candidate is generated by adding a Gaussian perturbation to the evaluated hyperparameter setting whose trajectory has contributed to the current front the most. It has been shown that such a candidate search guarantees asymptotic convergence to the global optimum [Regis and Shoemaker, 2007, Wang et al., 2023]. More importantly, it enables the simultaneous computation of TEHVI for multiple candidates in a batch to significantly improve efficiency. Please refer to Appendix B for more details.
5.2 Early Stopping and Augmentation
@@
$t^{*}=\sup\big{\{}t\in\mathbb{T}\mid\bm{\mu}(\bm{x}^{\prime},t)-\beta^{\frac{1}{2}}\bm{\Sigma}(\bm{x}^{\prime},t)\prec\bm{y},\exists\bm{y}\in F\big{\}},$
where $\beta$ is a predetermined constant controlling the confidence level. Inspired by the Lower Confidence Bound (LCB) [Srinivas et al., 2009], the conservative stopping epoch is the maximum number of epochs after which any future model performance is unlikely to improve the current front with a high probability. As the iterative learning procedure proceeds by epoch, the trajectory prediction is progressively updated based on the new observations on $Trj(\bm{x}^{\prime})$ and so is the conservative stopping epoch. The iterative learning procedure terminates once the current training epoch $t^{\prime}$ exceeds the conservative stopping epoch $t^{*}$.
Figure 2: Box plots for 20 problems derived from ZDT1, ZDT2, DTLZ1, DTLZ2, and DTLZ7. Each algorithm runs for 20 independent trials. The logarithm of Hypervolume difference is computed at the end of each trial and is normalized in $[0,1]$.
Consequently, at the end of each iteration of TMOBO, we obtain the temporary datasets $Z^{\prime}=\{(\bm{x}^{\prime},t)\}_{t=1}^{t^{\prime}}$ and $Y^{\prime}=\{\bm{y}(\bm{x}^{\prime},t)\}_{t=1}^{t^{\prime}}$, where the actual stopping epoch $t^{\prime}$ can take any value from $\{1,\dots,t_{max}\}$. However, retaining all new observations in $Z^{\prime}$ and $Y^{\prime}$ for training GP models in subsequent iterations is inefficient, especially when $t_{max}$ is large. Following the active data augmentation used in [Nguyen et al., 2020, Belakaria et al., 2023], we opt to augment only a subset of $Z^{\prime}$ to the primary dataset $Z$ by sequentially selecting an input $(\bm{x}^{\prime},t)\in Z^{\prime}$ at which the GP predictive uncertainty is highest. As more than one GP models are used to approximate the objectives, similar to the scenario considered in [Belakaria et al., 2023], the model uncertainty of $(\bm{x}^{\prime},t)$ is computed as the sum of normalized variances predicted by each GP model at $(\bm{x}^{\prime},t)$. Through this method, we can effectively control the increase in the size of the training set while ensuring the accuracy and training efficiency of GP models.
@@ 6 Numerical Experiments
In this section, we conduct a comprehensive empirical analysis of the performance of TMOBO on several synthetic and real-world benchmark problems that are formulated as EMOHPO. Recall that EMOHPO is essentially a multi-objective optimization problem defined over the combined space of hyperparameter settings and training epochs. Therefore, we compare TMOBO with several state-of-the-art MOBO methods, namely ParEGO [Knowles, 2006], $q$EHVI [Daulton et al., 2020], and $q$NEHVI [Daulton et al., 2021]. Notably, as discussed in Sections 1 and 3, we exclude multi-fidelity optimization algorithms from the comparison because they do not consider the possibility that earlier epochs (or lower fidelity levels) could contribute to Pareto-optimal set or front. As a result, these methods are not applicable to solving EMOHPO. To ensure a fair comparison, we further enhance selected MOBO methods by collecting all the observations $\{\bm{y}(\bm{x},1),\dots,\bm{y}(\bm{x},t)\}$ into their results when sampling at the pair $(\bm{x},t)$. For clarity, in the following discussion, we use ParEGO-T to represent the enhanced version of ParEGO, and similarly for $q$EHVI-T and $q$NEHVI-T.
Considering that all algorithms are stochastic, we perform 20 independent trial runs of each algorithm on each test problem. The logarithm of the HV difference between the front found by an algorithm and the true Pareto-optimal front is computed to measure the performance [Daulton et al., 2020, 2021]. Since the true Pareto-optimal front is generally unknown in practice, we approximate it by aggregating all observed solutions across all algorithms and trials and extracting the non-dominated set to form an empirical Pareto front. Although the empirical front may not fully capture the true Pareto front, this ensures an unbiased comparison, as no algorithm is favored a priori, and each method’s performance is evaluated relative to the best-known solutions discovered collectively. Similarly, the reference point is set as the least favorable solution with each dimension corresponding to the worst observed value of an objective so that it upper bounds all observations in the objective space for HV computation. Further details on algorithm configurations and corresponding sensitivity analyses are provided in Appendices C.1 and D, respectively. In the rest of this section, we showcase and analyze the main numerical results of this study but refer interested readers to Appendix E for detailed results and additional experiments.
6.1 Synthetic Simulations